# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Deep Learning

### v4

@@ 5.1 Experimental Setup
Figure 2: Left: Comparing performance of Bella models with their baseline (Base) counterparts on vision benchmarks. The results confirm Bella models are well-performing and on par with their baseline Ensemble, SVGD, and VBL whilst consuming only a fraction of cost (see Table 1 grey columns for corresponding costs). The best-performing results are in bold, the second-best in underline, and the least favorable are in italic for emphasis. Right: Bella achieves similar Accuracy with full SVGD (Base) only with a fraction of cost.
Neural Architecture. In our experiments, we employed the CLIP ViT-B/32 model (Ilharco et al. 2021), a pre-trained variant utilizing contrastive supervision from image-text pairs, as initially introduced in the seminal CLIP research (Radford et al. 2021). We conducted end-to-end fine-tuning of the image encoder, adjusting all model parameters, a strategy typically yielding higher accuracy than training only the final linear layer. For both our ensemble-based methods, we consider conventional ensemble and SVGD. We use the average logits (unnormalized outputs) to produce the output (Gontijo-Lopes, Dauphin, and Cubuk 2021). For VQA task, we employ the SoTA LLaVA-1.5-7B (Liu et al. 2023b) to showcase the effectiveness of Bella on large-scale network architecture. Details of the hyper-parameters are discussed in Appendix I.
@@ 5.2 Cost Efficiency
The accuracy comparison of our Bella models with their respective base models, as well as with Vision Bayesian Lora (VBL)—a derived Laplace’s approximation from (Yang et al. 2024) for vision tasks—is delineated in the left table in Figure 2. The respective costs comparison on training (in terms of trainable parameters) is shown in Table 1—notably, the grey columns are the ones used for generating the results in the left table in Figure 2.
                         Models Bseline Models (SVGD) Bella Models (SVGD) Single
$n=3$ $n=5$ $n=20$ $n=40$ $n=3$ $n=5$ $n=20$ $n=100$ $n=1$
Storage Consumption (MB) 1321 $2$ 8868 17735 436 439 460 572 433
Table 1: Computational cost to train different models based on CLIP architecture for different datasets. Notably, with SVGD Baseline Models, we can only train up to $n$=40 particles on a A6000 48 GB GPU, while we can increase to more than 100 parameter particles with our Bella method with negligible increase of GPU consumption. The grey columns correspond to costs (parameters) of models in Figure 2.
@@
Notably, in Figure 2 with Bella SVGD, with 1.6% of the trainable parameters in comparison to the Single baseline, leads to approximately 2% and 10% increase in performance on the OOD tasks of DomainNet and CAMELYON17, respectively; whilst achieving comparable performance with the current SVGD implementation (SVGD baseline model).
Along with Figure 2, Table 1 demonstrates the primary benefit of our approach—the significant reduction in memory and storage needs. For example, Table 1, we obtain approximately a 5$\times$ reduction in model size—that is from 2,222 MB for a 5 particle SVGD baseline model to just 440 MB for a Bella SVGD model with 5 particles. This efficiency not only reduces the demands on GPUs but also minimizes potential I/O bottlenecks.
Moreover, the reduced GPU demand, as shown in Table 1, facilitates larger mini-batch sizes during training to speedup the training process. More crucially, it enables the enhancement of the Bayesian posterior’s parameter particles to over 100, a feat previously unattainable with current SVGD implementations. Significantly, constructing a 100 parameter Bayesian approximation consumes only 5.19 GB memory compared to current SVGD implementations for Bayesian models (SVGD base) needing over 6 GB for even a merely 3 particle approximation.
Interestingly, our on-par results of Bella models with current SVGD and ensemble approximations of the posterior provide empirical evidence that with constrained model parameters, it is still possible to reach the diverse modes of the posterior. Further, the results support recent conjectures on mode connectivity (Gueta et al. 2023).
@@
Table 2: The out-of-distribution generalization performance of Bella, measured by accuracy (for number of particles $n=5$). Each column represents a specific shift, either real (DomainNet) or artificial (CIFAR-10-C). Further results can be found in Appendix D.
5.3 Out-Of-Distribution (OOD) Datasets
@@
First, we use the DomainNet dataset, which is one of the most diverse domain adaptation datasets and spans a wide range of visual styles, from real images to abstract art. This variety provides a challenging test bed for algorithms aiming to bridge different visual domains. Our study involves training the CLIP network on the ‘Real’ subset of DomainNet and evaluating its generalization across various domains. Additionally, we use CIFAR-10-C, a corrupted version of CIFAR-10, to further assess the generalization and robustness of models trained on CIFAR-10 datasets. The results for both datasets are presented in LABEL:tab:domainnet_cifar10-c.
Furthermore, the STL-10 dataset, which has significant label overlap with CIFAR-10, serves as a relevant OOD test case for CIFAR-10 models (see Appendix B).
Results from LABEL:tab:domainnet_cifar10-c and Appendix B show that Bella models, demonstrating significantly better efficiency, achieve competitive performance compared to more resource-intensive implementations with Ensemble and SVGD baselines across both DomainNet and CIFAR-10-C benchmarks. All Bayesian approximations, including our scalable and efficient method, outperform the single model baseline.
@@ 5.4 Comparing Uncertainty Estimations
Bayesian models capable of providing a theoretical basis for measuring model uncertainty. Also known as epistemic uncertainty, refers to uncertainty stemming from limitations in our knowledge or understanding of the underlying data generating process or the model itself. One of the ways to quantify model uncertainty is through mutual information estimates, following  (Gal et al. 2016).
@@
Results.  Figure 3 demonstrates the effectiveness of our approach to estimate uncertainty. Our Bella perform similarly to the SVGD base model, with a slightly better uncertainty on misclassified images of CAMELYON17 and corrupted CIFAR-10-C datasets (under brightness corruption with the maximum intensity), see details and other corruption types in Appendix C.
Figure 3: Evaluation of uncertainty estimations using Mutual Information on CAMELYON17 and CIFAR-10-C datasets. $\uparrow$ MI for Misclassified Examples is better—denoted by the distribution shifting $\rightarrow$. In contrast, $\downarrow$ MI for Correctly Classified Examples is better—denoted by the distribution shifting $\leftarrow$. Notably, Single base models are not able to capture epistemic uncertainty (Gal et al. 2016).
@@ 5.5 Robustness against Adversarial Examples
In this section, we examine the resilience of our proposed Bella against adversarial attacks, specifically employing the $L_{\infty}$ Fast Gradient Sign Method (FGSM) across various attack budgets as detailed in Figure 4. This analysis aims to benchmark the robustness of our method in comparison to traditional models under adversarial conditions. We employ the robustness benchmark (Papernot et al. 2018) to deploy the attack on CIFAR-10 test set and report results in Figure 4.
Figure 4: Comparison of model robustness to $L_{\infty}$ FGSM adversarial attacks across varied attack budgets. The evaluation is on the CIFAR-10 dataset. Bella achieves comparable performance to current, baseline SVGD and Ensemble (number of particles is equal to $5$) implementations whilst demonstrating significantly better robustness (higher accuracy) than a Single baseline model.
@@ 5.6 Ablation Studies
This section undertakes a series of ablation studies to examine the effects of various components within Bella. Our analysis includes an exploration of the training costs associated with different ranks and their consequent influence on model performance. Given that Bella incorporates multiple parameter particles, we also delve into how varying the number of these particles affects Bella’s efficacy. Additionally, we explore the application of low-rank adapters across different layers and assess their impact. Further details on other studies are in Appendix E. We show in Table 7 that we achieve state-of-the-art performance on CAMELYON17.
@@
To thoroughly assess the influence of rank, we conduct our analysis on the challenging large-scale CAMELYON17 dataset. The results, depicted in Figure 5 (left), elucidate the relationship between rank size and performance. While utilizing a smaller rank considerably reduces the parameter space, it also restricts Bella’s learning capacity, hindering its ability to achieve optimal performance. In contrast, increasing the rank to 4 significantly enhances Bella’s efficiency. However, performance tends to plateau at a rank of 16, indicating a saturation point.
@@
Datasets Metrics Bella Models Baseline Models (Base)
@@
Table 3: Calibration comparison of Bella models (SVGD and Ensemble) with baseline methods across multiple datasets. The evaluation considers calibration metrics (ECE, MCE) and discriminative capability (Brier Score, AUROC). Lower values for ECE, MCE, and Brier Score, along with higher values for AUROC, indicate better performance.
@@
The plot in Figure 2 (right) for the CAMELYON17 dataset begins with Bella at $r=2$ (comprising $5$ particles) and concludes with full SVGD (SVGD base) using the same number of particles. As $r$ increases, the accuracy of Bellas improves, eventually plateauing at approximately 95.08%, which is comparable to the accuracy of full SVGD (95.21%). This highlights the efficiency and advantages of our proposed Bella. Remarkably, Bella achieves performance on par with full SVGD models despite utilizing a significantly smaller pool of trainable parameters—approximately 0.3% for $r=4$—compared to the more parameter-intensive alternatives.
@@
Calibration Study. Table 3 presents a comparison of the Bella models (SVGD and ensemble) against baseline models, including Variational Inference (VI) (Kim and Hospedales 2023) and Stochastic-Gradient Langevin Dynamic (SGLD) (Welling and Teh 2011), using several key metrics: Expected Calibration Error (ECE), Maximum Calibration Error (MCE), Brier score, and Area Under the Receiver Operating Characteristic curve (AUROC). Lower ECE and MCE values indicate that the model’s probability estimates are more reliable. A lower Brier score reflects better accuracy in the model’s probability predictions, while a higher AUROC demonstrates superior discrimination between classes. Across the datasets, the Bella models demonstrate performance comparable to the baseline models. This suggests that the Bella models, while utilizing far fewer parameters for training, offer a strong alternative to traditional baseline approaches while maintaining competitive calibration and Discriminative capability.
5.7 Visual Question Answering
@@
In this section, we extend the application of our Bella to another challenging vision task, Visual Question Answering (VQA), as detailed in (Antol et al. 2015). We leverage the state-of-the-art, pre-trained, large multi-modal model LlaVA  (Liu et al. 2023b, a) for this purpose. Utilizing LlaVA transforms VQA into a process where an image and a natural-language question are inputs, and the model generates a free-form, open-ended text answer. Answering questions in VQA requires various intelligent capabilities, including not only image recognition but also complex reasoning. For this task, we employ VQA v2 (Antol et al. 2015) dataset containing 204,721 images, more than 1 Billion (1B) questions and 10B ground-truth answers in total. There are three main types of answers: Yes/No, a Number, and Other.
@@
Model. In our experiments, we employed our proposed Bella on top of LlaVA-1.5-7B (Liu et al. 2023b, a). Further details about the dataset, model and metrics are deferred to Appendix H.
@@
Accuracy.  The primary outcomes for Yes/No and Number answer queries are detailed in Table 4, respectively. We chose these question types to ensure a fair comparison and avoid semantic mismatches in open-ended answers in Others. A distinctive feature of Bella is its reduced uncertainty estimates for currect predictions, coupled with increased uncertainty estimates for incorrect ones. Moreover, it surpasses the Single base model regarding Accuracy and Exact Match metrics.
@@
Notably, our method is efficient, particularly in contrast to the resource-intesive baseline Bayesian models tailored for this task, e.g. Abbasnejad et al. (2019, 2020b, 2020a). The computational requirements for using these baselines render the application of full SVGD or ensemble alternatives impractical, thereby highlighting the impact on practical applications by harnessing the effectiveness of Bella.
@@
Models Ent. Corrects ($\downarrow$) Ent. Incorrect ($\uparrow$) Accuracy ($\uparrow$) Match ($\uparrow$)
@@
Table 4: LLaVA-VQA Results: evaluating the accuracy and entropy of correct vs incorrect predictions.
Model Uncertainty and Human Confidence. Further, we investigate the relationship between model uncertainty attained from Bella and human confidence, facilitated by multiple annotators in the VQA dataset. For this, we measure the correlation between the model disagreement by measuring the predictive entropy vs the human annotations of the answers.
(a) Yes/No questions.
(b) Number questions.
Figure 6: Correlation between model certainty and human confidence. This negative correlation suggests that the model’s entropy can serve as a reliable gauge of certainty, mirroring human judgment in its response to varying levels of uncertainty.
In Figure 6 we shows the correlation between the entropy of the model outputs and human confidence levels. By setting a specific threshold for the model’s entropy, we effectively bifurcate our predictions into two distinct categories: those with lower entropy fall into the ‘Low Entropy’ group, signalling reduced uncertainty within the model’s assessments, while predictions surpassing this entropy threshold are allocated to the ‘High Entropy’ segment, indicative of greater uncertainty. Intriguingly, our observations reveal a consistent negative correlation between the model’s entropy levels and human confidence across different types of queries, such as Yes/No and Number questions. This pattern suggests that the model’s entropy can serve as a reliable gauge of certainty, mirroring human judgment in its response to varying levels of uncertainty.
@@ 6 Conclusion
In this paper, we introduce an innovative approach for creating an efficient Bayesian Neural Network (BNN) approximation using only a base pre-trained model. Our approach, called Bella, demonstrates remarkable compatibility with full-rank BNN approximations like SVGD or Ensemble methods, outperforming single-network solutions across a range of tasks, from pure classification to out-of-distribution (OOD) generalization and uncertainty quantification. Our research paves the way for effective and efficient BNNs, facilitating the development of reliable and robust machine-learning models. Bella reinforces the effectiveness of employing such a simple, efficient and effective training method with diverse representations over the conventional practice of single fine-tuning.

### v5

@@ 5.1 Experimental Set-up
Table 2: Comparing Bella models with their baseline (Base) counterparts on vision benchmarks. Bella models are well-performing and on par with their baseline Ensemble, SVGD, and VBL whilst consuming only a fraction of cost. The best-performing results are in bold, the second-best in underline, and the least favorable are in italic for emphasis.
Neural Architecture. In our experiments, we employed the CLIP ViT-B/32 model [47], a pre-trained variant utilizing contrastive supervision from image-text pairs, as initially introduced in the seminal CLIP research [48]. We conducted end-to-end fine-tuning of the image encoder, adjusting all model parameters, a strategy typically yielding higher accuracy than training only the final linear layer. For both our ensemble-based methods, we consider conventional ensemble and SVGD. We use the average logits (unnormalized outputs) to produce the output [49]. For VQA task, we employ the SoTA LLaVA-1.5-7B [18] to showcase the effectiveness of Bella on large-scale network architecture. Detailed hyper-parameters are in the Appendix.
@@ 5.2 Cost Efficiency
Cost comparisons (in terms of trainable parameters, memory and storage) are shown in Table 1. The performance comparison of our Bella models with their respective base models, as well as with Vision Bayesian Lora (VBL)—a derived Laplace’s approximation from [24] for vision tasks—is shown in Table 2. In particular, the gray columns Table 1 indicate the models used to generate the results in Table 2.
@@
Notably, in Table 2 with Bella SVGD, with 1.6% of the trainable parameters in comparison to the Single baseline, leads to approximately 2% and 10% increase in performance on the OOD tasks of CAMELYON17 and DomainNet, respectively; whilst achieving comparable performance with the current SVGD implementation (SVGD baseline model).
Along with Table 2, Table 1 demonstrates the primary benefit of our approach—the significant reduction in memory and storage needs. For example, as seen in Table 1, we obtain approximately a 5$\times$ reduction in model size—that is from 2,222 MB for a 5 particle SVGD baseline model to just 439 MB for a Bella SVGD model with 5 particles. This efficiency not only reduces the demands on GPUs but also minimizes potential I/O bottlenecks.
Moreover, the reduced GPU demand, as shown in Table 1, facilitates larger mini-batch sizes during training to speedup the training process. More crucially, it enables one to enhance the Bayesian posterior’s parameter particles to over 100, a feat previously unattainable with current SVGD implementations. Significantly, constructing a 100 parameter Bayesian approximation consumes only 5.19 GB memory compared to current SVGD implementations for Bayesian models (SVGD base) needing over 6 GB for even a mere 3 particle approximation.
Interestingly, our results for Bella models are on-par with SVGD and ensemble approximations of the posterior. This provides empirical evidence that with constrained model parameters, it is still possible to reach the diverse modes of the posterior. Further, the results support recent conjectures on mode connectivity [50].
@@
Table 3: The out-of-distribution generalization performance of Bella, measured by accuracy (for number of particles $n=5$). Each column represents a specific shift, either real (DomainNet) or artificial (CIFAR-10-C).
5.3 Out-Of-Disrtibution (OOD) Datasets
@@
First, we use the DomainNet dataset, which is one of the most diverse domain adaptation datasets and spans a wide range of visual styles, from real images to abstract art. This variety provides a challenging test bed for algorithms aiming to bridge different visual domains. Our study involves training the CLIP network on the ‘Real’ subset of DomainNet and evaluating its generalization across various domains. Additionally, we use CIFAR-10-C, a corrupted version of CIFAR-10, to further assess the generalization and robustness of models trained on CIFAR-10 datasets. The results for both datasets are presented in Table 3.
Furthermore, the STL-10 dataset, which has significant label overlap with CIFAR-10, serves as a relevant OOD test case for CIFAR-10 models (see the Appendix).
Results from Table 3 and the Appendix show that Bella models, demonstrating significantly better efficiency, achieve competitive performance compared to more resource-intensive implementations with Ensemble and SVGD baselines across both DomainNet and CIFAR-10-C benchmarks. All Bayesian approximations, including our scalable and efficient method, outperform the single model baseline.
@@ 5.4 Comparing Uncertainty Estimations
Bayesian models capable of providing a theoretical basis for measuring model uncertainty. Also known as epistemic uncertainty, refers to uncertainty stemming from limitations in our knowledge or understanding of the underlying data generating process or the model itself. One of the ways to quantify model uncertainty is through mutual information estimates, following  [51].
@@
Results.  Figure 2 demonstrates the effectiveness of our approach to estimate uncertainty. Our Bella perform similarly to the SVGD base model, with a slightly better uncertainty on misclassified images of CAMELYON17 and corrupted CIFAR-10-C datasets (under brightness corruption with the maximum intensity), see details in the Appendix.
Figure 2: Evaluation of uncertainty estimations using Mutual Information on CAMELYON17 and CIFAR-10-C datasets. $\uparrow$ MI for Misclassified Examples is better—denoted by the distribution shifting $\rightarrow$. In contrast, $\downarrow$ MI for Correctly Classified Examples is better—denoted by the distribution shifting $\leftarrow$.
@@ 5.5 Robustness against Adversarial Examples
In this section, we examine the resilience of our proposed Bella against adversarial attacks, specifically employing the $L_{\infty}$ Fast Gradient Sign Method (FGSM) across various attack budgets as detailed in Figure 3. This analysis aims to benchmark the robustness of our method in comparison to traditional models under adversarial conditions. We employ the robustness benchmark [52] to deploy the attack on CIFAR-10 test set and report results in Figure 3.
Figure 3: Comparison of model robustness to $L_{\infty}$ FGSM adversarial attacks across varied attack budgets on the CIFAR-10 dataset.
@@ 5.6 Ablation Studies
This section undertakes a series of ablation studies to examine the effects of various components within Bella. Our analysis includes an exploration of the training costs associated with different ranks and their consequent influence on model performance. Given that Bella incorporates multiple parameter particles, we also delve into how varying the number of these particles affects Bella’s efficacy. Additionally, we explore the application of low-rank adapters across different layers and assess their impact. Further details on other studies are in the Appendix. We show in the Appendix that we achieve state-of-the-art performance on CAMELYON17.
@@
To assess the influence of rank, we conduct an analysis on the challenging, large-scale, CAMELYON17 task. The results, reported in Figure 4 (left), elucidate the relationship between rank size and performance. While utilizing a smaller rank considerably reduces the parameter space, it also restricts Bella’s learning capacity, hindering its ability to achieve optimal performance. In contrast, increasing the rank to 4 significantly enhances performance. However, performance tends to plateau at a rank of 16, indicating a saturation point.
@@
Figure 5: Bella achieves similar Accuracy with full SVGD (Base) only with a fraction of cost.
Data. Metric Bella Base
@@
Table 4: Calibration comparison of Bella models (SVGD and Ensemble) with baseline methods across multiple datasets.
@@
The plot in Figure 5 for the CAMELYON17 dataset begins with Bella at $r=2$ (comprising $5$ particles) and concludes with full SVGD (SVGD base) using the same number of particles. As $r$ increases, the accuracy of Bellas improves, eventually plateauing at approximately 95.08%, which is comparable to the accuracy of full SVGD (95.21%). This highlights the efficiency and advantages of our proposed Bella. Remarkably, Bella achieves performance on par with full SVGD models despite utilizing a significantly smaller pool of trainable parameters—approximately 0.3% for $r=4$—compared to the more parameter-intensive alternatives.
@@
Calibration Study. Table 4 presents a comparison of the Bella models (SVGD and ensemble) against baseline models, including Variational Inference (VI) [53] and Stochastic-Gradient Langevin Dynamic (SGLD) [38], using several key metrics: Expected Calibration Error (ECE), Maximum Calibration Error (MCE), Brier score, and Area Under the Receiver Operating Characteristic curve (AUROC). Lower ECE and MCE values indicate that the model’s probability estimates are more reliable. A lower Brier score reflects better accuracy in the model’s probability predictions, while a higher AUROC demonstrates superior discrimination between classes. Across the datasets, the Bella models demonstrate performance comparable to the baseline models. This suggests that the Bella models, while utilizing far fewer parameters for training, offer a strong alternative to traditional baseline approaches while maintaining competitive calibration and Discriminative capability.
5.7 Generalization to a Visual Question Answer
@@
In this section, we extend the application of our Bella to another challenging vision task, Visual Question Answering (VQA), as detailed in [54]. We leverage the state-of-the-art, pre-trained, large multi-modal model LlaVA  [18, 55] for this purpose. Utilizing LlaVA transforms VQA into a process where an image and a natural-language question are inputs, and the model generates a free-form, open-ended text answer. Answering questions in VQA requires various intelligence capabilities, including not only image recognition but also complex reasoning. For this task, we employ VQA v2 [54] dataset containing 204,721 images, more than 1 Billion (1B) questions and 10B ground-truth answers in total. There are three main types of answers: Yes/No, a Number, and Other.
@@
Model. We employed our proposed Bella on top of LlaVA-1.5-7B [18, 55]. Further details about the dataset, model and metrics are deferred to the Appendix.
@@
Accuracy.  The primary outcomes for Yes/No and Number answer queries are detailed in Table 5. We chose these question types to ensure a fair comparison and avoid semantic mismatches in open-ended answers in Others. A distinctive feature of Bella is its reduced uncertainty estimates for correct predictions, coupled with increased uncertainty estimates for incorrect ones. Moreover, it surpasses the Single base model regarding Accuracy and Exact Match metrics.
@@
Notably, our method is efficient, particularly in contrast to the resource-intesive baseline Bayesian models tailored for this task (e.g. [56, 57, 13]). The computational requirements for using these baselines render the application of full SVGD or ensemble alternatives impractical, thereby highlighting the impact on practical applications by harnessing the effectiveness of Bella.
@@
Models Ent. Corrects ($\downarrow$) Ent. Incorrect ($\uparrow$) Acc. ($\uparrow$) Match ($\uparrow$)
@@
Table 5: LLaVA-VQA Results: evaluating the accuracy and entropy of correct vs. incorrect predictions (Ens: Ensemble).
Model Uncertainty and Human Confidence. Further, we investigate the relationship between model uncertainty attained from Bella and human confidence, facilitated by multiple annotators in the VQA dataset. For this, we measure the model disagreement by measuring the correlation between predictive entropy and human annotations of the answers.
Figure 6: Correlation between model certainty and human confidence. This negative correlation suggests that the model’s entropy can serve as a reliable gauge of certainty, mirroring human judgment.
In Figure 6 we show the correlation between the entropy of the model outputs and human confidence levels. By setting a specific threshold for the model’s entropy, we effectively bifurcate our predictions into two distinct categories: those with lower entropy fall into the ‘Low Entropy’ group, signaling reduced uncertainty within the model’s assessments, while predictions exceeding the threshold are allocated to the ‘High Entropy’ group, indicative of greater uncertainty. Intriguingly, our observations reveal a consistent negative correlation between the model’s entropy levels and human confidence across the different query types (Yes/No and Number questions). This pattern suggests that the model’s entropy can serve as a reliable gauge of certainty, mirroring human judgment in its response to varying levels of uncertainty.
@@ 6 Conclusion
We present Bella, an innovative efficient Bayesian Neural Network (BNN) approximation using a base pre-trained model. Bella achieves remarkable compatibility with full-rank BNN methods like SVGD and Ensembles, surpassing single-network solutions in classification, OOD generalization, and uncertainty quantification. This approach enables scalable, reliable BNN implementations, demonstrating the benefits of simple and efficient training over traditional single fine-tuning techniques. Code released at https://bnn-bella.github.io/BNN-Bella/